OpenStudy (anonymous):

After the softball game, Alise's coach bought each of the 13 players a double-scoop or triple-scoop ice cream cone at the local ice cream shop. The coach bought a total of 33 scoops of ice cream. Use a model to determine how many girls got a triple-scoop ice cream cone.

OpenStudy (tad1):

x + y = 13 2x + 3y =33 mult 1st eq by -2 -2x -2y =-26 2x +3y = 33 add the two equations y = 7 so 7 girls got 3 scoops and 6 girls got 2 scoops. check: 7 check 7*3 = 21 2*6 = 12 21 + 12 = 33 it checks

OpenStudy (anonymous):

Thanks! :)